OpenStudy (anonymous):

Given the two expressions shown below: square root of 3 plus square root of 2 square root of 5 plus square root of 6 Which statement best describes the two expressions? Both are rational. Both are irrational. A is rational, but B is irrational. A is irrational, but B is rational. I think it's b.

OpenStudy (anonymous):

@bohotness

OpenStudy (bohotness):

yes

OpenStudy (anonymous):

Okay, can you check another answer for me?

OpenStudy (bohotness):

yes

OpenStudy (anonymous):

The following function represents the value of a house, in dollars, after x years: f(x) = 242,000(1.04)x What does 242,000 represent? The present value of the house The value of the house after x years The increase in the value of the house each year The increase in the value of the house after x years A?

OpenStudy (danjs):

Is 1.04 raised to the x power?

OpenStudy (anonymous):

Yes.

OpenStudy (danjs):

I just know, i think it should be an exponential function

OpenStudy (danjs):

when x = 0 y = 242000(1.04)^0 a number to the zero power = 1 y(0) = 24000

OpenStudy (anonymous):

So is the answer A.. or?

OpenStudy (danjs):

the 242000 represents the initial value of the house at time zero

OpenStudy (danjs):

so i guess A would be it, if the present is considered time t=0,, then yes it is the present value of the house

OpenStudy (anonymous):

Okay, thank you! Can you help me with another? Sorry, cramming segment exams. :/